Let E and F be two events of an experiment with sample space S. Suppose  P(E)= 0.4,...

  1. Let E and F be two events of an experiment with sample space S. Suppose  P(E)= 0.4, P(F)=0.3, P(E  U F) =0.5,  Find  P(F|E)   and determine if the two events are independent.

    A) P(F|E)= 3/4, E and F are independent.

    B) P(F|E)= 3/4, E and F are not independent.

    C) P(F|E)=1/2 , E and F are independent.

    D) P(F|E)= 1/2, E and F are not independent.
Answer #1

Given, P(E)=0.42 PIF) 0.3, PLEUF) = 0.5 PLEUF) = PIE) +PLF) - PENE) 0.5 = 0.4 +0.3 - PLENE > PIENF) = 0.7 -0.5 PLEND = 0.2 PL